Given that the point (-40, -9) is on the terminal side of an angle, θ, find the exact, simplified value of the following:
sin=
cos=
tan=
csc=
cot=
sec=

Found 3 solutions by Mariam8775, ikleyn, Alan3354:
Answer by Mariam8775(1)

Sin= -12.7
Cos= 167.3
Tan= 12.7

Answer by ikleyn(52787)

.
Mariam, you are HEAVILY wrong, demonstrating absolute absence knowledge on the subject.


Answer by Alan3354(69443)

Given that the point (-40,-9) is on the terminal side of an angle, θ, find the exact:
-----------------
x = -40
y = -9
---
r = sqrt(x^2+y^2) = 41
======================
sin = y/r
cos = x/r
tan = y/x
---
cot = 1/tan
sec = 1/cos
csc = 1/sin